# NRSeg: Noise-Resilient Learning for BEV Semantic Segmentation via Driving World Models

### v1

@@ III-D The Overall Loss
where $Loss_{gt}$ is the Dice loss with the PGCM module for the source domain. $Loss_{pl}$ and $Loss_{e}^{t}$ are the supervision of the pseudo-probabilistic labels of the target domain, constrained by the L2 loss. $\lambda_{1}=0.5$ and $\lambda_{2}=0.1$ are referenced from the works [21, 48]. The $\lambda_{3}$ is set as $1$. $\beta$ is controlled by a sigmoid rampup function, which starts at 00 and gradually increases to $0.1$ when the training round is halfway through.
@@ IV-B Implementation Details
Training Settings: The unsupervised domain adaptation and semi-supervised learning frameworks in this section are both established on the Mean Teacher (MT) architecture [49]. The learning momentum of the teacher model is $\alpha=0.99$.
LSS [6] is chosen as the pipeline of the BEV semantic segmentation task, where the size of a perspective image is $(128,352)$. The BEV semantic segmentation includes $6$ static classes. The perception range of BEV space is $(-50,50)m$ and the solution is $0.5m$, which is similar to PCT [3]. We use the mean Intersection over Union (mIoU) as the main evaluation metric. When batch is $12$, the learning rate is set to $3e^{-3}$. PolyLR optimizer is used for all tasks with a minimum learning rate of $1e^{-5}$. All experiments are conducted on NVIDIA RTX A6000 GPUs.
@@ IV-C Performance Comparison
Table VI: Analysis on the amount of synthetic data on UDA. ‘O-S’ means the original size of the labeled data, whereas ‘N-S’ represents the added labeled data.
@@
1.8k 17.3
0.6k 21.0
0.6k 20.7
0.6k 20.0
1.2k 21.7
1.8k 22.6
2.4k 22.0
IV-D Ablation Studies
Ablation on the PGCM Module: To verify the broad applicability of the proposed PGCM module, this subsection assesses its effectiveness across synthetic data generated by different world models. As shown in Table IV, the PGCM module demonstrates consistent effectiveness in improving BEV semantic segmentation accuracy, whether using synthetic data from individual models or combined data from multiple models. Furthermore, we conducted comparative experiments with alternative loss optimization approaches. Specifically, following the approach in the work [50], we implemented two distinct optimization schemes.
[…]
The primary objective of M1 is to employ a score $S$ to control the predicted probability $P_{+}$, preventing it from easily converging to $1$ during learning. In contrast, M2 utilizes $S$ to accelerate and guide the synthetic data to align with the labeled data direction. The baseline of this section is built on the SSL of $1/8$ setting. As shown in Table V, the results demonstrate that the former has a minimal impact on learning, whereas the latter hurts model performance. This negative effect stems from the overly efficient convergence of synthetic data, which leads to erroneous learning directions. Our approach successfully rectifies these suboptimal learning directions and reaches a better performance.
Ablation on the Amount of Training Data: This paper aims to utilize diversified synthetic data to enhance the capabilities of BEV semantic segmentation models. In this subsection, we examine the impact on the amount of diversified synthetic data on BEV model learning. As shown in Table VI, when the amount of newly added synthetic data is less than the original labeled data, the improvement in accuracy is proportional to the amount of data. However, when the amount of synthetic data exceeds that of the original labeled data, the accuracy begins to decline. This occurs because the model tends to learn from the majority of the data, where the noisy synthetic data can misguide the learning. Moreover, due to variations in image generation quality across different world models, it can be observed that the slightly inferior BEVControl [13] provides only limited improvement $2.7\%$ mIoU. The more performant controlled PerlDiff [1] contributes a $3.7\%$ mIoU performance improvement.
Ablation on Synthetic Image Quality: To further analyze the quality of synthetic data generation, we compared the learning performance between synthetic and real images in semi-supervised tasks. Here, we divide the validation set into two parts, one part serves as newly labeled training data, while the other part continues to function as the validation set. As shown in Table VII, incorporating synthetic images as training data yielded inferior results compared to real images, indicating certain generative errors in the quality of synthetic data. However, after optimizing the learning direction with the proposed PGCM module, the accuracy is improved significantly, demonstrating the effectiveness of PGCM.
Table VII: Ablation on the synthetic image quality. ‘Ori.’ represents the original image data. ‘Perl.’ denotes the synthetic data of the world model PerlDiff [1].
@@ IV-D Ablation Studies
Ori. 0.3k 0.2k 31.6
Perl. 0.3k 0.2k 30.8
Perl. 0.3k 0.2k 31.4
Table VIII: Ablation on different text conditions.
@@
Table IX: Ablation on the method of perspective mask generation. ‘Sys. Mask’ denotes the method to process synthetic data, whereas ‘Ref. Mask’ denotes the reference mask. ‘GTProj’ projects the BEV GT mask to the perspective view.
Sys. Mask Ref. Mask mIoU
Baseline 21.7
@@
Ablation on Different Text Settings: The world model can generate synthetic data under diverse environments by modifying text inputs, which proves highly beneficial for cross-domain adaptation learning. In our study of day-night cross-domain adaptation, we systematically compare the effects of different text configurations on domain transfer performance. As quantitatively demonstrated in Table VIII, compared to official text containing mixed weather conditions, synthetic data generated exclusively under nighttime settings further improves the accuracy by $1.2\%$ in mIoU.
@@
Ablation on the Perspective Mask Method: The perspective consistency score is dependent on the accuracy of mask generation for synthetic data. Therefore, this section presents a comprehensive analysis of different mask models. As shown in Table IX, we compare the performance between Mask2Former [45] and SAN [46] models, with the former demonstrating superior results in our study. The baseline is the UDA of $Boston\xrightarrow{}Singapore$ setting with the synthetic data from PerlDiff and MagicDrive. The Mask2Former model is trained on Cityscapes, a dataset specifically designed for autonomous driving scenarios, which is similar to the nuScenes dataset. Thus, the perspective pseudo-labels generated on nuScenes data exhibit higher accuracy, leading to more reliable perspective geometry consistency scores. Furthermore, we considered another scenario where the original real images corresponding to the synthetic data exist. In this case, reference masks can also be generated using real images and a semantic segmentation model. The results show comparable performance to that obtained using BEV mask projection.
@@ IV-E Visualization Analyses
Figure 4: Visualization results and consistent scores of synthetic data from PerlDiff [1] and MagicDrive [2]. (a) The first column displays the reference image of the dataset and the back-projected road mask. The latter two columns present the generated synthetic data and predicted masks from Mask2Former [45]. (b) Consistency scores R of the synthetic data corresponding to the image examples from top to bottom.
@@ IV-F New-Split Fully-Supervised Performance
The previous StreamMapNet [21] on vectorized high definition mapping introduced a novel dataset splitting strategy for nuScenes, specifically designed to evaluate BEV model generalization. Compared to the original dataset splitting, the newly defined training and validation sets exhibit a significant distribution gap, leading to degraded generalization performance in BEV models. To verify the effect of synthetic data on model generalization, we evaluate the proposed noise-resilient learning framework under fully supervised tasks on the newly split dataset. As demonstrated in Table X, the proposed NRSeg significantly enhances the accuracy of BEV semantic segmentation with an improvement of $3.3\%$ in mIoU, proving that incorporating diversely distributed data effectively improves the generalization capability.
Table X: Results on the new-split nuScenes. It is implemented for the full-supervised BEV semantic segmentation.
@@
Table XI: Results on cross-dataset domain adaptation. ‘VT’ denotes the view transformer module.
@@
IV-G Cross-Dataset Adaptation Performance
Due to the strong dependency of BEV models on training data characteristics, domain adaptation across datasets remains a significant challenge owing to inconsistencies in sensor configurations and data collection environments. Therefore, this section provides an analysis of our method for cross-dataset domain adaptation between Argoverse [51] and nuScenes. Our experiments use Inverse Perspective Mapping (IPM) as the view transformation module, an approach that has been proven more suitable for cross-dataset BEV mapping [24]. As shown in Table XI, after sequentially adding synthetic data from PerlDiff and MagicDrive, the cross-dataset BEV domain adaptation capability shows significant improvement. The inclusion of synthetic data from a single world model yielded a $0.5\%$ precision improvement, while combining synthetic data from two world models delivered a $2.7\%$ precision boost. This result further validates the effectiveness of our proposed noise-resilient learning framework for cross-dataset domain adaptation tasks.
@@ V Conclusion
The current research in this work incorporates synthetic data as an augmented data distribution into the source domain for learning. Whether synthetic data from driving world models can be directly used for learning in the context of source-free domain adaptation remains a question worthy of investigation. We will further investigate this critical question and dedicate efforts to develop a principled solution.

### v2

@@ III-D The Overall Loss
where $Loss_{gt}$ is the Dice loss with the PGCM module for the source domain. $Loss_{pl}$ and $Loss_{e}^{t}$ are the supervision of the pseudo-probabilistic labels of the target domain, constrained by the L2 loss. $\lambda_{1}=0.5$ and $\lambda_{2}=0.1$ are referenced from the works [49, 18]. The $\lambda_{3}$ is set as $1$. $\beta$ is controlled by a sigmoid ramp-up function, which starts at 0 and gradually increases to $0.1$ when the training round is halfway through.
@@ IV-B Implementation Details
Training Settings: The unsupervised domain adaptation and semi-supervised learning frameworks in this section are both established on the Mean Teacher (MT) architecture [35]. Here, the student model is jointly trained on source and target domain data with strong data augmentation. The strong data augmentation methods include resize, rotation, cropping, flipping, and colorjitter. The teacher model is updated via momentum based on the weights of the student model. The learning momentum of the teacher model is $\alpha=0.99$.
LSS [29] is chosen as the pipeline of the BEV semantic segmentation task, where the size of a perspective image is $(128,352)$. The BEV semantic segmentation includes $6$ static classes. The perception range of BEV space is (50,50)m(-50,50)m and the solution is $0.5m$, which is similar to PCT [14]. We use the mean Intersection over Union (mIoU) as the main evaluation metric. When batch is $12$, the learning rate is set to $3e^{-3}$. PolyLR optimizer is used for all tasks with a minimum learning rate of $1e^{-5}$. All experiments are conducted on NVIDIA RTX A6000 GPUs.
@@ IV-C Performance Comparison
IV-D Ablation Studies
Figure 3: Analysis of the convergence of training loss for the PGCM module. ‘w’ indicates that the PGCM module is used, and ‘w/o’ means not used.
Ablation on Denoising Effect of PGCM Module: To verify the broad applicability of the proposed PGCM module, this subsection assesses its denoising effectiveness across synthetic data generated by different world models. The result is shown in Table IV. As observed from the FID scores, the synthetic data generated by PerfDiff [50] exhibits relatively high quality. Consequently, in the cross-domain BEV segmentation task, the synthetic data from this model contributes the most significant positive impact, achieving $21.0\%$ in mIoU. Additionally, the PGCM module demonstrates consistent effectiveness in improving BEV semantic segmentation accuracy, whether using synthetic data from individual models or combined data from multiple models.
In addition, this section analyzes the convergence after loss optimization. As shown in Fig. 3, this section analyzes the loss convergence under different cross-domain settings. With the inclusion of the PGCM module, the model still demonstrates efficient convergence and achieves higher accuracy, with an increase of $1.2\%$ mIoU and $0.4\%$ mIoU, respectively.
@@ IV-D Ablation Studies
Table VI: Results of different loss functions. The synthetic data from the PerfDiff [50] is selected as the baseline.
Loss Function Sys. Data PGCM mIoU
Focal Loss Perl. 18.7
DICE Loss Perl. 21.0
Perl. 21.5
Table VII: Analysis on the loss weight of different distributions. $Loss_{gt}$ corresponds to the multinomial distribution loss, and $loss_{e}$ corresponds to the Dirichlet distribution loss.
$Loss_{gt}$ $Loss_{e}$ PGCM mIoU
1 0.1 23.7
1 0.5 24.0
1 1.0 24.6
Ablation on the Structure of PGCM Module: This section first conducted comparative experiments with alternative loss optimization approaches. Specifically, following the work [20], we implemented two distinct optimization schemes.
$\text{M1}=1-\frac{\textbf{R}*2*P_{+}}{\textbf{R}*(P_{+}+P_{-})+N_{+}+\sigma}.$ (31)
$\text{M2}=1-\frac{(2-\textbf{R})*2*P_{+}}{(2-\textbf{R})*(P_{+}+P_{-})+N_{+}+\sigma}.$ (32)
The primary objective of M1 is to employ a score $S$ to control the predicted probability $P_{+}$, preventing it from easily converging to $1$ during learning. In contrast, M2 utilizes $S$ to accelerate and guide the synthetic data to align with the labeled data direction. The baseline of this section is built on the SSL of $1/8$ setting. As shown in Table VI, the results demonstrate that the former has a minimal impact on learning, whereas the latter hurts model performance. This negative effect stems from the overly efficient convergence of synthetic data, which leads to erroneous learning directions. Our approach successfully rectifies these suboptimal learning directions and reaches a better performance.
Subsequently, in this section, we further conduct experiments targeting different loss functions, where the Focal Loss is selected as the comparative loss function. The result is shown in Table VI. It can be observed that the performance of Focal Loss is inferior to that of DICE loss when synthetic data is added to the source domain data. This is because the Focal Loss strongly constrains learning at the individual pixel level, which can lead to erroneous fitting. In contrast, DICE loss exhibits higher tolerance for subtle structural misalignments.
Ablation on the Loss Weight of BiDPP: Both distributions in the BiDPP module are subject to label constraints. This section further analyzes the impact of the constraint loss weights. As shown in Table VII, it can be observed that when the weight of the multinomial distribution loss $Loss_{gt}$ is fixed at $1$, the performance is optimal when the weights of the Dirichlet distribution loss $Loss_{e}$ are relatively balanced. This configuration allows both to function synergistically.
Table VIII: Analysis on the amount of synthetic data on UDA. ‘O-S’ means the original size of the labeled data, whereas ‘N-S’ represents the added labeled data.
@@
18k 17.3
6k 21.0
6k 20.7
6k 20.0
12k 21.7
18k 22.6
24k 22.0
Ablation on the Amount of Training Data: This paper aims to utilize diversified synthetic data to enhance the capabilities of BEV semantic segmentation models. In this subsection, we examine the impact on the amount of diversified synthetic data on BEV model learning. As shown in Table VIII, when the amount of newly added synthetic data is less than the original labeled data, the improvement in accuracy is proportional to the amount of data. However, when the amount of synthetic data exceeds that of the original labeled data, the accuracy begins to decline. This occurs because the model tends to learn from the majority of the data, where the noisy synthetic data can misguide the learning. Moreover, due to variations in image generation quality across different world models, it can be observed that the slightly inferior BEVControl [45] provides only a limited improvement of $2.7\%$ in mIoU. The more performant controlled PerlDiff [50] contributes a $3.7\%$ mIoU performance improvement.
Ablation on Synthetic Image Quality: To further analyze the quality of synthetic data generation, we compared the learning performance between synthetic and real images in semi-supervised tasks. Here, we divide the validation set into two parts; one part serves as newly labeled training data, while the other part continues to function as the validation set. As shown in Table IX, incorporating synthetic images as training data yielded inferior results compared to real images, indicating certain generative errors in the quality of synthetic data. However, after optimizing the learning direction with the proposed PGCM module, the accuracy is improved significantly, demonstrating the effectiveness of PGCM.
Table IX: Ablation on the synthetic image quality. ‘Ori.’ represents the original image data. ‘Perl.’ denotes the synthetic data of the world model PerlDiff [50].
@@
Ori. 3k 2k 31.6
Perl. 3k 2k 30.8
Perl. 3k 2k 31.4
Ablation on Different Text Settings: The world model can generate synthetic data under diverse environments by modifying text inputs, which proves highly beneficial for cross-domain adaptation learning. In our study of day-night cross-domain adaptation, we systematically compare the effects of different text configurations on domain transfer performance. As quantitatively demonstrated in Table X, compared to official text containing mixed weather conditions, synthetic data generated exclusively under nighttime settings further improves the accuracy by $1.2\%$ in mIoU.
Table X: Ablation on different text conditions.
@@
Table XI: Ablation on the method of perspective mask generation. ‘Sys. Mask’ denotes the method to process synthetic data, whereas ‘Ref. Mask’ denotes the reference mask. ‘GTProj’ projects the BEV GT mask to the perspective view.
Sys. Mask Ref. Mask PGCM mIoU
21.7
@@
Ablation on the Perspective Mask Method: The perspective consistency score is dependent on the accuracy of mask generation for synthetic data. Therefore, this section presents a comprehensive analysis of different mask models. As shown in Table XI, we compare the performance between Mask2Former [4] and SAN [42] models. The baseline is the UDA of $Boston\xrightarrow{}Singapore$ setting with the synthetic data from PerlDiff and MagicDrive, where PGCM does not work. Overall, the consistency scores derived from the masks generated by both pre-trained models can effectively mitigate noise issues and contribute to accuracy improvements. Additionally, the Mask2Former model is trained on Cityscapes, a dataset specifically designed for autonomous driving scenarios, which is similar to the nuScenes dataset. The SAN model focuses on generalization capabilities under open-vocabulary conditions. Therefore, our analysis indicates that the pre-trained Mask2Former model, which focuses on traffic data, generates more precise pseudo-labels. Furthermore, we considered another scenario where the original real images corresponding to the synthetic data exist. In this case, reference masks can also be generated using real images and a semantic segmentation model. The results show comparable performance to that obtained using BEV mask projection.
Figure 5: Visualization results and consistent scores of synthetic data from PerlDiff [50] and MagicDrive [9]. (a) The first column displays the reference image of the dataset and the back-projected road mask. The latter two columns present the generated synthetic data and predicted masks from Mask2Former [4]. (b) Consistency scores R of the synthetic data corresponding to the image examples from top to bottom. (c) Compared to the reference image, the synthetic data within the red bounding box exhibits a slight offset in the road structure. Accordingly, the drivable area follows changed.
@@ IV-E Visualization Analyses
This paper designs a framework at both the data and architectural levels to prevent noisy synthetic data from overfitting to misaligned label data. As shown in Fig. 5(c), there are discrepancies in the road structure between the synthetic data and the reference image. The proposed method guides the model to learn the correct regions from the synthetic data.
@@ IV-F New-Split Fully-Supervised Performance
The previous StreamMapNet [49] on vectorized high definition mapping introduced a novel dataset splitting strategy for nuScenes, specifically designed to evaluate BEV model generalization. Compared to the original dataset splitting, the newly defined training and validation sets exhibit a significant distribution gap, leading to degraded generalization performance in BEV models. To verify the effect of synthetic data on model generalization, we evaluate the proposed noise-resilient learning framework under fully supervised tasks on the newly split dataset. As demonstrated in Table XII, the proposed NRSeg significantly enhances the accuracy of BEV semantic segmentation with an improvement of $3.3\%$ in mIoU, proving that incorporating diversely distributed data effectively improves the generalization capability.
IV-G Cross-Dataset Adaptation Performance
Due to the strong dependency of BEV models on training data characteristics, domain adaptation across datasets remains a significant challenge owing to inconsistencies in sensor configurations and data collection environments. Therefore, this section provides an analysis of our method for cross-dataset domain adaptation between Argoverse [39] and nuScenes. Our experiments use Inverse Perspective Mapping (IPM) as the view transformation module, an approach that has been proven more suitable for cross-dataset BEV mapping [19]. As shown in Table XIII, after sequentially adding synthetic data from PerlDiff and MagicDrive, the cross-dataset BEV domain adaptation capability shows significant improvement. The inclusion of synthetic data from a single world model yielded a $0.5\%$ precision improvement, while combining synthetic data from two world models delivered a $2.7\%$ precision boost. This result further validates the effectiveness of our proposed noise-resilient learning framework for cross-dataset domain adaptation tasks.
Table XII: Results on the new-split nuScenes. It is implemented for the full-supervised BEV semantic segmentation.
@@ IV-G Cross-Dataset Adaptation Performance
Table XIII: Results on cross-dataset domain adaptation. ‘VT’ denotes the view transformer module.
@@ V Conclusion
Limitations: The proposed denoising framework effectively mitigates the inherent noise in such synthetic data. However, this framework requires co-training on the original source-domain data. When the volume of available source domain data is limited, the utility of noise synthetic data remains constrained.
Future Work: In scenarios where only a pre-trained model on the source domain is available, approaches such as meta-learning can be employed to learn common task features from noisy synthetic data, thereby guiding learning on the unlabeled target domain.